The length of a rectangle is 1 cm longer than its width. If the diagonal of the rectangle is 4 cm, what are the dimensions (the length and the width) of the rectangle?
Answer by stanbon(48516)

The length of a rectangle is 1 cm longer than its width. If the diagonal of the rectangle is 4 cm, what are the dimensions (the length and the width) of the rectangle?
-----------------
Let the width be x; the length will be "x+1".
-----------
Using Pythagoras:
diagonal^2 = x^2 + (x+1)^2
16= 2x^2 + 2x +1
2x^2+2x-15 = 0
x = [-2+-sqrt(124)]/4
Positive answer:
x = [-2+sqrt(124)]/4
x = 2.28388... (width)
x+1 = 3.28388... (length)
================
Cheers,
Stan H.